OpenStudy (anonymous):

Which of the following sums would be under the radical symbol to find the distance between the points (7, -1) and (-8, -9)? 1^2+ 10^2 1^2+ 8^2 15^2+ 8^2

OpenStudy (radar):

Determine the x distance and the y distance 7-(-8)=15 -1-(-9)=8 take it from there.